A bike is traveling at 7.4 m/s when it slows at a constant rate and comes to a stop after 3.4 seconds what is the bicycles acceleration?
1
Expert's answer
2020-10-12T07:46:57-0400

By definition, the acceleration is:


"a = \dfrac{v_{final} - v_{initial}}{t}"

where "v_{final} = 0m\/s" is the final velocity of the bike (since it comes to rest), "v_{initial} = 7.4m\/s" is the intial velocity of the bike and "t = 3.4s" is the time of motion. Thus, have:

"a = \dfrac{v_{final} - v_{initial}}{t} = \dfrac{0m/s -7.4m/s}{3.4s} \approx 2.18m/s^2"

Answer. 2.18 m/s^2.